5. In the circuit below, calculate the output volatge (v.) in time and phasor domains. (30 points) 60 S2 cos(t – 15° 10 mF

need ans asap
Answer #1

Sol given that fon 15 H 3 Vo ☺vi CIOMFF Vi = 60364–150) = 54-15° volt w=l rad/sec XL = LW=5*1=51 xe = tw = 1080x1053 xx=1001No G3s shoo + o) vo (-360 +30.6 +1) = V; Vo = Eje 1300$+1) No Citilius Vo = Colt-150) .44L-84.98 No=0:08763(+-15°+84.989) Vo=